Find the perimeter and area of the sector 1c 셋
Answer #1

In the diagram the radius is 4 and \theta = 1 radian

The perimeter = 2r + Length of arc = r\theta = (2*4) + (4*1) = 12

Area of the sector = (0.5) * r2 * \theta = 0.5 * 16 * 1 = 8